Membership The SIPB is a group of undergraduates, graduate students, and other members of the MIT community sharing an interest in computer science and in community service. Not all members are computer scientists; some are involved in physics, math and a variety of other fields. SIPB members work in various labs and departments around campus, providing an informal network of contacts between independent groups of computer users.

The SIPB has two kinds of members: full and associate. A full member of the SIPB must be a student currently registered for courses at the Institute. All officers of the Board must be full members, and only full members may vote.

All members of the Board are expected to attend Board meetings and to contribute towards the activities of the Board on a regular basis. This includes staffing the SIPB office, participating in policy and design discussions, and helping out with the constant stream of jobs associated with providing services to the community.

Below is a list of the 152 SIPB members active during the 1982-1992 academic years, and their period of activity within the 1982-1992 time frame:

Harold Ancell (1982 - 1988)

Derek Atkins (1989 - 1992)

Richard Barbalace (1992)

Michael Barrow (1989 - 1991)

Elisa Barzelay (1982 - 1983)

Richard Basch (1987 - 1992)

Michael Bauer (1991 - 1992)

Alan Bawden (1982 - 1983)

Don Becker (1983 - 1985)

Sharon Belville (1988 - 1989)

Jay Berkenbilt (1988 - 1991)

Pandora Berman (1982 - 1983, 1992)

Jeff Bigler (1990 - 1992)

Deborah Birnby (1991)

Julie Bourne (1989 - 1992)